# One Queue Is All You Need: Resolving Head-of-Line Blocking in Large Language Model Serving

### v1

@@ 7 Plan Generator
$ct_{g,j}$ Request group completion time
@@
Overall Modeling Approach. The goal of the stochastic programming solver is to find an assignment of request groups to the virtual queues so that all SLOs are met. To model SLO attainment, we define a penalty term for each request group, which is the difference between completion time and SLO value. If SLOs are met, all penalty terms would be smaller than 0. Given the SLOs as the inputs to the stochastic programming model, we obtain the request group completion time estimation from the RCT estimator (see section 6) to estimate the defined penalty terms. The completion time for a request group is the sum of the waiting time for request groups ahead in the virtual queue (from Equation 2), the completion time for the request group (from Equation 5), and swap times associated with transferring model weights into GPU memory.  [^2] We measure swap times from model load time profiling. Note that effects associated with hardware and model heterogeneity (such as token throughput and eviction vs. swap) that impact request group completion time are captured by the RCT estimator profiling.
Mathematical Definitions of Constraints. Now, we describe each of the constraints mathematically in further detail. We assume that each virtual queue can have a maximum length, and every request group is assigned to one of the positions in the virtual queue. Equation 6 models request group assignment to a position in the virtual queue.
$\sum_{g}\sum_{j}x_{g,i,j}=1\forall i\quad\sum_{i}x_{g,i,j}=1\forall g,j$ (6)
@@
$m_{g,j}=\sum_{i}\text{models}_{i}\times x_{g,i,j}\forall g,j$ (7)
$slo_{g,j}=\sum_{i}\text{slos}_{i}\times x_{g,i,j}\forall g,j$ (8)
The transition between two different models is captured in Equation 9. While inequalities cannot be directly modeled as constraints, we apply the standard big-M method to reduce it further [7].
$t_{g,j}=(m_{g,j-1}\neq m_{g,j})\forall g,j$ (9)
The cumulative completion times of all positions in the virtual queue would be the sum of waiting time, completion times, and swap times as represented in Equation 10. We assume a constant conservative estimate of swap time, but this can be further improved in future work.
$\text{ct}_{g,j}=\sum_{i}\sum_{k}^{j-1}W_{g,i}\times x_{g,i,k}+\sum_{k}^{j-1}t_{g,k}\times S+\sum_{i}C_{g,i}\times x_{g,i,j}\forall g,j$ (10)
Note that completion times are a stochastic variable as group completion times $C_{g,i}$ are represented with a probability distribution (i.e., $C$ from Equation 5).
The penalty for completion times would simply be the difference between the completion time and the SLO value, as shown in Equation 11.
$p_{g,j}=ct_{g,j}-slo_{g,j}\forall g,j$ (11)
The final constraint is that all penalty values should be less than 0 (i.e., all SLOs are satisfied) with a high probability.
$P(p_{g,j}\leq 0)>\Delta\forall g,j$ (12)
Optimization Goal. The stochastic programming model aims to minimize the total penalty for SLO violations.
$\min(\sum_{g}\sum_{j}E[p_{g,j}])$ (13)
8 Evaluation
@@
    QLM performance with respect to SLO attainment and request throughput in multi-model serving (section 8.1) and single-model serving (section 8.2),
@@
  3. (c)
    Accuracy of the RCT estimator in request completion time estimation on production LLM serving traces,
@@
Experiment Setup. We evaluate QLM on multiple varying-sized open-source LLMs: Mistral-7B [19], Vicuna-13B [5], and Llama-70B [50]. Our test cluster consists of 80–100 GPUs of two types: NVIDIA A10 (24 GB memory) and NVIDIA A100 (80 GB memory). The setup represents both model and device heterogeneity. To evaluate the benefit of QLM, we consider the following three baseline mechanisms: (1) EDF(Earliest Deadline First): Requests are sorted by their SLO values such that requests with the smallest SLO values are at the front of the virtual queue, (2) vLLM[21]: Requests use the default first-come-first-serve (FCFS) scheduler in vLLM, and (3) SHEPHERD[60]: Requests use dynamic batching and an ILP formulation for ordering and placement. Note that SHEPHERD cannot be easily extended to work with continuous batching as the LP formulation assumes fixed batches with deterministic execution times.
Workloads. We consider the following four workloads that are derived from the requirements of a production cloud service provider: [$\mathbf{W_{A}}$] (Multi-Model)Requests to different models arrive at the queue, and the number of served models is larger than the number of corresponding LLM serving instances. Consequently, more than one model needs to be multiplexed on the same LLM serving instance. Additionally, based on our insights from production request patterns, we assume that $>50\%$ of models are “tail models”, i.e., they are not used often and only occasionally need to be swapped in place of the frequently used models. Each request type has varied SLO requirements based on the downstream application requirements. For example, latency-sensitive services have a completion time SLO <30s, and batch jobs have SLOs in the order of several minutes. While our choice of SLO values is motivated by production requirements, we also provide robustness analysis for alternative SLO values in Appendix A.1. [$\mathbf{W_{B}}$] (Single-Model)Requests to a single model arrive at the request queue, and no model swapping is required. Similar to $W_{A}$, these requests have multiple SLOs dependent on the underlying application. [$\mathbf{W_{C}}$] (MegaPrompt)The request workload consists of several “mega prompts” in addition to the workload from $W_{A}$. These mega prompts have a large number of input and output tokens that occupy a large percentage of GPU memory and potentially prevent requests in the waiting queue from executing.
Figure 8: Multi-model request serving throughput .
Figure 9: Multi-model SLO satisfaction.
Figure 10: Multi-model LSO ablation study.
8.1 Multi-Model Evaluation
We run workload $W_{A}$ to evaluate the multi-model LLM serving performance on A100 GPUs with respect to request throughput and end-to-end latency SLO satisfaction. Additionally, we also provide an ablation study to understand the contribution of each LSO to the overall QLM performance.
Request Throughput and SLO Attainment. Figure 10 shows the request throughput (i.e., requests served per second) for $W_{A}$ comparing QLM with the baseline mechanisms for varying percentage of tail model arrival rates (normalized to peak tail model demand). QLM provides up to 3–4× higher throughput due to the following factors: (1) The use of request groups minimizes repeated swapping required as the model would only be swapped in once per request group instead of per individual request, and (2) The plan generator couples every tail model with another frequently accessed model to minimize swaps while maintaining an equal distribution of queue sizes. The improvement in request throughput directly maximizes the percentage of SLO satisfied for all requests. Figure 10 shows the percentage of SLO satisfied for the latency-sensitive services against the percentage of their occurrence in the whole serving workload. When these latency-sensitive services constitute less than 5% of the request queue, QLM satisfies more than 90% of all SLO values. As the relative percentage of latency-sensitive service requests increases, no combination of requests would be able to meet all SLOs, and the plan generator would fail to return a solution. In such a scenario, a scale-up action is required to add more GPU devices. We perform this scale-up action to enable 100% SLO attainment if the current GPU capacity is insufficient. The baselines perform worse compared to QLM because none of them consider the impact of model swapping. Other limitations of the baselines are discussed in Section 8.2.
Contribution of Each LSO. Each of the five LSOs used by QLM, including request pulling/eviction, GPU-CPU state swapping, model warm start, and load balancing, contributes to either the latency and/or the throughput of the serving system. Figure 10 shows the impact of removing each LSO on QLM performance for $W_{A}$. The model warm start LSO contributes the most to QLM performance for both SLOs and throughput, as multiple models need to be multiplexed on the same LLM serving instance. Additionally, the other LSOs contribute primarily to the latency SLO attainment.
@@ 8.1 Multi-Model Evaluation
Figure 11: Single model request serving throughput.
@@
Figure 12: Single model SLO satisfaction.
@@
Figure 13: Single model LSO ablation study.
8.2 Single-Model Evaluation
@@
We run workload $W_{B}$ on A100 GPUs to evaluate the single-model LLM serving performance regarding the request throughput, SLO attainment, and LSO contribution ablation study (similar to the multi-model evaluation in section 8.1).
@@
Request Throughput and SLO Attainment. Figure 13 shows the percentage of SLOs that are satisfied by QLM and the baseline systems. Similar to the multi-model case, we find that when the queue primarily consists of latency-sensitive services, none of the systems can satisfy the SLOs. This is because the minimum serving time is much longer than the specified SLO. As the number of latency-sensitive service requests decreases, QLM performs the best in satisfying the maximum number of SLOs. Specifically, it performs better than the baseline mechanisms because: (a) Compared to vLLM, QLM is able to move latency-sensitive service requests ahead in the queue, (b) Compared to EDF, QLM enables appropriate selection between GPU-CPU state swapping and request eviction LSOs, and (c) Compared to SHEPHERD, QLM uses continuous batching as opposed to static batch size and models the auto-regressive LLM nature with the RCT estimator to increase request throughput. We find that the advantages of QLM with respect to smart selection among various LSOs, continuous batching, and appropriate request prioritization help with improving request throughput. Figure 13 shows the request throughput for QLM and the individual baseline mechanisms. QLM achieves higher throughput, i.e., 20% higher compared to vLLM and EDF, and 50% higher than SHEPHERD.
@@
Contribution of Each LSO. Figure 13 shows the impact of removing each LSO considered by the backend LLM serving instance in QLM. Scheduling and request eviction contribute significantly to latency reduction for latency-sensitive services and consequently increase the number of SLOs met. GPU-CPU state swap increases request throughput by swapping the KV cache into CPU memory. Finally, model swapping has no impact on this workload as a single model is being served.
Figure 14: Impact of hardware heterogeneity.
Figure 15: Impact of mega prompt arrivals.
Figure 16: Impact of increasing queue size on SLO satisfaction.
8.3 QLM Robustness Analysis
@@
Hardware Heterogeneity. We run $W_{A}$ on a mix of A10 and A100 GPUs to evaluate the robustness of QLM performance in heterogeneous hardware setup. Figure 16 shows request throughput when the cluster has varying ratios of A10 to A100 GPUs. The A10 is a lower-end GPU with $\sim$3× lower GPU memory and thus is only capable of serving a much lower request throughput compared to the A100 GPU. QLM takes into account this difference between request throughput across GPUs via the RCT estimator with offline profiling, and the plan generator proportionally assigns a lower number of requests to the A10 GPU compared to the A100. On the other hand, if we use a round-robin policy for request assignment to the LLM serving instances (while using default QLM policy per instance), the load would be distributed equally, leading to higher queue drain times for the A10 GPU. Additionally, we also observe that the benefit of QLM is more compared to a random policy when the heterogeneity of the cluster is higher. When the A10 GPUs constitute 20–50% of the cluster (more heterogeneous), the improvement of QLM over random policy is 2–5× higher compared to a 100% A10 or 100% A100 composition (more homogeneous).
@@
Mega Prompt Workload. The RCT estimator of QLM takes into account input and output token distribution when estimating the request completion time. Consequently, when there are distinct token distributions, such as in workload setup $W_{C}$, QLM is able to load balance them intelligently across LLM serving instances to minimize the queue drain time. For example, in workload $W_{C}$, the “mega prompts” use a large number of tokens, and their KV cache occupies the entire GPU memory, causing head-of-the-line blocking for the regular requests in the queue. The optimal policy, as identified by QLM, in such a scenario would be to allocate all the regular requests to another LLM serving instance. Note that request eviction is not an option if all SLOs are tight. Figure 16 shows the benefit of QLM for workload $W_{C}$. The relative benefit of QLM is highest for a few mega prompts because the regular requests can be moved to another GPU. As the percentage of mega prompts increases, there is no option but to assign them to different LLM serving instances, causing inevitable HOL blocking, and the benefit of QLM reduces. In such a case, we would need to perform a scale-up action and add more GPU devices to the cluster to continue maintaining SLOs.
@@
Figure 17: Accuracy of RCT estimator.
Figure 18: Impact of request group size on QLM performance.
Figure 19: QLM Overhead.
RCT Estimator Accuracy. The RCT estimator calculates the request completion time based on initial profiling of the model and hardware setup. This initial profiling time is negligible as only a single batch of requests need to be run on the GPU. As described in Section 6, QLM does not assume that the exact output tokens are known ahead of time, but instead uses the workload output token distribution. Figure 19 shows the estimated request completion time vs. the actual request completion time for the RCT estimator across different models and hardware configurations. Overall, we find that the RCT estimator has a high accuracy in estimating request completion times with an $R^{2}$ (coefficient of determination) value of 0.99 (out of 1.0). While the RCT estimator is highly accurate in estimating request completion time, it is not perfect. There could be requests with an outlier number of output tokens, leading to underestimation and potential SLO violations. However, current LLM serving systems typically have a hard limit on the maximum number of output tokens [21, 36], which eliminates the presence of these outliers.
@@
Impact of Request Group Size. QLM sets the request group size as a multiple ($\delta$) of the average batch size. The exact $\delta$ value depends on the acceptable trade-off between the overhead of running the plan generator and the granularity of decision-making. As $\delta$ becomes smaller, QLM achieves a finer granularity of decision-making, leading to improved performance. However, the overhead leads to delayed decision-making. Figure 19 demonstrates this tradeoff between performance degradation (caused by changing granularity in decision making) and overhead of the plan generator when varying $\delta$. At $\delta=16$, the overhead is smallest, but decision-making granularity is coarse, leading to sub-optimal decisions (such as imbalance between virtual queue sizes of LLM serving instances). In contrast, at $\delta=1$, the performance degradation is minimal, but overhead is much higher. We choose $\delta=4$, as it results in nearly zero performance degradation, compared to $\delta=1$, while maintaining a low overhead.
@@
Scalability and Overhead. The overhead of QLM largely depends on the time required to solve the stochastic programming formulation required by the plan generation. In Figure 19, we show the time required to solve for the plan generator with varying queue sizes in terms of the number of requests. As the basic unit of the solver is a single request group, the model and GPU configurations with a larger request group size would be able to handle a much larger queue size for the same overhead. Consequently, configurations with a large request group size, such as an A100 with a 7B model, can handle a maximum queue size of 400K requests at a 5s overhead per request group (i.e., 5 ms per request).
9 Related Work
General ML Model-Serving Systems. Traditional model-serving systems provide functionalities such as scheduling, placement, batching, and autoscaling. Clipper [9], TensorFlow-Serving [35], MArk [59], InferLine [8], Shepherd [60], and Clockwork [16] are some earlier work on serving traditional ML models like ResNet that are relatively small. INFaaS [41] and Cocktail [17] propose a model-less serving framework to automate the model selection and autoscaling to meet SLOs. However, they fail to consider the autoregressive property of LLMs. On the other hand, advanced autoscaling techniques are complementary to QLM.

### v2

@@ 7. Global Scheduler
$wt_{g,j}$ Request group waiting time
@@
Overall Modeling Approach. The goal of the linear programming solver is to find an assignment of request groups to the virtual queues so that all SLOs are met. To model SLO attainment, we define a penalty term for each request group, which is the difference between waiting time and SLO value. If SLOs are met, all penalty terms would be smaller than 0. Given the SLOs as the inputs to the linear programming model, we obtain the request group waiting time estimation from the RWT estimator to estimate the defined penalty terms. The worst-case waiting time for a request group is the sum of the waiting time for request groups ahead in the virtual queue from the same model (from Equation 2), the completion time for the request group for different models (from Equation 5), and swap times associated with transferring model weights into GPU memory  [^3] We measure swap times from model load time profiling.. Note that effects associated with hardware and model heterogeneity (such as token throughput and eviction vs. swap) that impact request group completion time are captured by the RWT estimator profiling.
Definitions of Constraints. Now, we describe each of the constraints in further detail. We assume that each virtual queue can have a maximum length, and every request group is assigned to one of the positions in the virtual queue. Equation 6 models request group assignment to a position in the virtual queue.
(6) $\sum_{g}\sum_{j}x_{g,i,j}=1\forall i\quad\sum_{i}x_{g,i,j}=1\forall g,j$
@@
(7) $m_{g,j}=\sum_{i}\text{models}_{i}\times x_{g,i,j}\forall g,j$
(8) $slo_{g,j}=\sum_{i}\text{slos}_{i}\times x_{g,i,j}\forall g,j$
The transition between two different models is captured in Equation 9. While inequalities cannot be directly modeled as constraints, we apply the standard big-M method to reduce the inequality into linear constraints (Cococcioni and Fiaschi, 2021).
(9) $t_{g,j}=(m_{g,j-1}\neq m_{g,j})\forall g,j$
The cumulative waiting times of all positions in the virtual queue would be the sum of waiting time, completion times, and swap times as represented in Equation 10. $C$ is the completion time of a request group from Equation 5.
(10) $\text{wt}_{g,j}=\sum_{i}\sum_{k}^{j-1}W_{g,i}\times x_{g,i,k}+\sum_{k}^{j-1}t_{g,k}\times S+\sum_{i}\sum_{k}^{j-1}C_{g,i}\times t_{g,k}\times x_{g,i,j}\forall g,j$
The penalty would simply be the difference between the waiting time and the SLO value, as shown in Equation 11.
(11) $p_{g,j}=wt_{g,j}-slo_{g,j}\forall g,j$
The final constraint is that all penalty values should be less than 0 i.e., all SLOs are satisfied.
(12) $p_{g,j}\leq 0\forall g,j$
Optimization Goal. The linear programming model aims to minimize the total penalty for SLO violations.
(13) $\min(\sum_{g}\sum_{j}p_{g,j})$
8. Evaluation
@@
    QLM performance with respect to SLO attainment and request throughput in single-model serving (section 8.1),
@@
    QLM performance with respect to SLO attainment and request throughput in multi-model serving (section 8.2),
  3. (c)
@@
    Accuracy of the RWT estimator in request waiting time estimation,
  5. (e)
@@
Experiment Setup. We evaluate QLM on multiple varying-sized open-source LLMs: Mistral-7B (Jiang et al., 2023), Vicuna-13B (Chiang et al., 2023), and Llama-70B (Touvron et al., 2023). We evaluate on a test bed consisting of GPUs of two types: 30 NVIDIA A10 (24 GB memory) and 50 NVIDIA A100 (80 GB memory). The setup represents both model and device heterogeneity. To evaluate the benefit of QLM, we consider the following three baseline mechanisms: (1) EDF(Earliest Deadline First): Requests are sorted by their SLO values such that requests with the smallest SLO values are at the front of the virtual queue, (2) vLLM(Kwon et al., 2023): Requests use the default first-come-first-serve (FCFS) scheduler in vLLM, and (3) SHEPHERD(Zhang et al., 2023b): Requests are served with dynamic batching and an ILP formulation for ordering and placement. Note that SHEPHERD cannot be easily extended to work with continuous batching as the LP formulation assumes fixed batches with deterministic execution times.
Figure 8. Distribution of input and output tokens in the shareGPT dataset.
[…]
Distribution of input and output tokens in the shareGPT dataset.
Workloads. We create our experimental workloads from the requirements of a production cloud service provider except for request arrival rates due to confidentiality reasons. Request arrivals are modeled with a Poisson distribution and queues are created by varying the arrival rates. Each workload trace uses 3,500 requests from the ShareGPT (team, 2024) dataset with input/output token distribution as shown in Figure 8. We classify all requests into three categories and define their SLO values accordingly: (1) Interactive:20s, (2) Batch-1:1 min, and (3) Batch-2:1 hour. Note that these SLOs are defined with respect to the 99th percentile value of the time to first token (TTFT). We test the experimental workload in the following three scenarios: [$\mathbf{W_{A}}$] Single-Model Interactive and Batch Workloadwhich consists of Batch-1, Batch-2, and Interactive requests for a single model, and no model swapping is required. [$\mathbf{W_{B}}$] Multi-Model Batch Workloadwhich consists of Batch-1 and Batch-2 requests. Batch-1 requests use two models: fine-tuned versions of Mistral 7B and Llama 70B. Batch-2 requests use three models: fine tuned versions of Vicuna 13B and Llama 70B. [$\mathbf{W_{C}}$] Single-Model MegaPrompt WorkloadThis request workload consists of several “mega prompts” in addition to the workload from $W_{B}$. To generate the mega prompt workload, we randomly select requests with total input and output tokens in the 3K – 4K range. These mega prompts have a large number of input and output tokens that occupy a large percentage of GPU memory and cause further HOL blocking.
We do not present results for the multi-model interactive workload because QLM assigns a separate GPU for each model, effectively reducing the workload to a single model workload ($W_{A}$). QLM’ global scheduler decides not to swap the models as the model swapping time exceeds the interactive request SLO (20s).
@@
Figure 9. Single model request serving throughput at 0.5K requests/s interactive arrival rate. Increased throughput corresponds to 1.1-2.3× GPU requirement reduction.
[…]
Single model request serving throughput at 0.5K requests/s interactive arrival rate. Increased throughput corresponds to 1.1-2.3× GPU requirement reduction.
@@
Figure 10. Single model SLO satisfaction for varying interactive request arrival rates for Vicuna 13B.
[…]
Single model SLO satisfaction for varying interactive request arrival rates for Vicuna 13B.
@@
Figure 11. Single model LSO ablation study at 0.5K requests/s interactive arrival rate for Vicuna 13B.
[…]
Single model LSO ablation study at 0.5K requests/s interactive arrival rate for Vicuna 13B.
8.1. Single-Model Evaluation
@@
We run workload $W_{A}$ on 50 A100 GPUs to evaluate the single-model LLM serving performance regarding the request throughput, SLO attainment, and LSO contribution ablation study (similar to the multi-model evaluation in section 8.2).
@@
Request Throughput and SLO Attainment. Figure 11 shows the percentage of SLOs that are satisfied by QLM and the baseline systems. We find that when the request arrival rate significantly exceeds the serving capacity, none of the systems can satisfy the SLOs. This is because the minimum serving time is much longer than the specified SLO. As the arrival rate of interactive requests decreases, QLM performs the best in satisfying the maximum number of SLOs. Specifically, it performs better than the baseline mechanisms because: (a) Compared to vLLM, QLM is able to move interactive service requests ahead in the queue, (b) Compared to EDF, QLM enables appropriate eviction of batch requests from the running queue, and (c) Compared to SHEPHERD, QLM uses continuous batching as opposed to static batch size and models the auto-regressive LLM nature with the RWT estimator to increase request throughput.
We find that the advantages of QLM with respect to smart selection among various LSOs, continuous batching, and appropriate request prioritization help with improving request throughput. Figure 11 shows the request throughput for QLM and the individual baseline mechanisms at arrival rate of 0.5K requests/s where QLM is able to achieve the maximum SLO satisfaction. QLM achieves higher throughput, i.e., 20% higher compared to vLLM and EDF, and 50% higher than SHEPHERD.
@@
Contribution of Each LSO. Figure 11 shows the impact of removing each LSO considered by the backend LLM serving instance in QLM. Request pulling contribute significantly to latency reduction for interactive services and consequently increase the number of SLOs met. Request eviction increases request throughput by swapping the KV cache into CPU memory. Finally, model swapping has no impact on this workload as a single model is being served.
Figure 12. Multi-model request serving throughput for varying Batch-1 request arrival rates. Increased throughput corresponds to 2-5× GPU requirement reduction.
[…]
Multi-model request serving throughput for varying Batch-1 request arrival rates. Increased throughput corresponds to 2-5× GPU requirement reduction.
Figure 13. Multi-model SLO satisfaction for varying Batch-1 request arrival rates.
[…]
Multi-model SLO satisfaction for varying Batch-1 request arrival rates.
Figure 14. Multi-model LSO ablation study for 0.25K requests/sec Batch-1 arrival rate.
[…]
Multi-model LSO ablation study for 0.25K requests/sec Batch-1 arrival rate.
8.2. Multi-Model Evaluation
We run workload $W_{B}$ to evaluate the multi-model performance on 50 A100 GPUs with respect to request throughput and SLO satisfaction. Additionally, we also provide an ablation study to understand the contribution of each LSO to the overall QLM performance.
Request Throughput and SLO Attainment. Figure 14 shows the request throughput (i.e., requests served per second) for $W_{B}$ comparing QLM with the baseline mechanisms for varying Batch-1 arrival rates. QLM provides up to 3–4× higher throughput due to the following factors: (1) The use of request groups minimizes repeated swapping required as the model would only be swapped in once per request group instead of per individual request, and (2) The global scheduler couples every Batch-1 model with another Batch-2 model to minimize swaps while maintaining an equal distribution of queue sizes.
The improvement in request throughput directly maximizes the percentage of SLO satisfied for all requests. Figure 14 shows the percentage of SLO satisfied for the interactive services against against their arrival rate. When the Batch-1 requests arrive at less than 0.5K requests/s, QLM satisfies more than 90% of all SLO values. As the arrival rate of Batch-1 service requests increases, no combination of requests would be able to meet all SLOs, and the global scheduler would fail to return a solution. In such a scenario, a scale-up action is required to add more GPU devices. We perform this scale-up action to enable 100% SLO attainment if the current GPU capacity is insufficient. The baselines perform worse compared to QLM because none of them consider the impact of model swapping. Other limitations of the baselines are discussed in Section 8.1.
Contribution of Each LSO. Each of the four LSOs used by QLM, including request pulling, request eviction, model swapping, and load balancing, contributes to either the latency and/or the throughput of the serving system. Figure 14 shows the impact of removing each LSO on QLM performance for $W_{B}$. The model warm start LSO contributes the most to QLM performance for both SLOs and throughput, as multiple models need to be multiplexed on the same LLM serving instance. Additionally, the other LSOs contribute primarily to the latency SLO attainment.
8.3. QLM Robustness Analysis
@@
Hardware Heterogeneity. We run $W_{A}$ on a mix of A10 and A100 GPUs to evaluate the robustness of QLM performance in heterogeneous hardware setup. Figure 17 shows request throughput when the cluster has varying ratios of A10 to A100 GPUs. The A10 is a lower-end GPU with $\sim$3× lower GPU memory and thus is only capable of serving a much lower request throughput compared to the A100 GPU. QLM takes into account this difference between request throughput across GPUs via the RWT estimator with offline profiling, and the global scheduler proportionally assigns a lower number of requests to the A10 GPU compared to the A100. On the other hand, if we use a round-robin policy for request assignment to the LLM serving instances (while using default QLM policy per instance), the load would be distributed equally, leading to higher queue drain times for the A10 GPU. Additionally, we also observe that the benefit of QLM is more compared to a random policy when the heterogeneity of the cluster is higher. When the A10 GPUs constitute 20–50% of the cluster (more heterogeneous), the improvement of QLM over random policy is 2–5× higher compared to a 100% A10 or 100% A100 composition (more homogeneous). Note that for experimental purposes, we increase the total number of A100s proportionately to demonstrate the impact of GPU imbalance.
@@
Mega Prompt Workload. The RWT estimator of QLM takes into account input and output token distribution when estimating the request waiting time. Consequently, when there are distinct token distributions, such as in workload setup $W_{C}$, QLM is able to load balance them intelligently across LLM serving instances to minimize the queue drain time. For example, in workload $W_{C}$, the “mega prompts” use a large number of tokens, and their KV cache occupies the entire GPU memory, causing head-of-the-line blocking for the regular requests in the queue. The optimal policy, as identified by QLM, in such a scenario would be to allocate all the regular requests to another LLM serving instance. Note that request eviction is not an option if all SLOs are tight. Figure 17 shows the benefit of QLM for workload $W_{C}$. The relative benefit of QLM is highest for a few mega prompts because the regular requests can be moved to another GPU. As the percentage of mega prompts increases, there is no option but to assign them to different LLM serving instances, causing inevitable HOL blocking, and the benefit of QLM reduces. In such a case, we would need to perform a scale-up action and add more GPU devices to the cluster to continue maintaining SLOs.
Figure 15. Impact of hardware heterogeneity.
[…]
Impact of hardware heterogeneity.
Figure 16. Impact of mega prompt arrivals.
[…]
Impact of mega prompt arrivals.
Figure 17. Impact of increasing queue size on SLO satisfaction.
[…]
Impact of increasing queue size on SLO satisfaction.
@@
Figure 18. Accuracy of RWT estimator.
[…]
Accuracy of RWT estimator.
Figure 19. Impact of request group size on QLM performance.
[…]
Impact of request group size on QLM performance.
Figure 20. QLM Overhead.
[…]
QLM Overhead.
RWT Estimator Accuracy. The RWT estimator calculates the request waiting time based on initial profiling of the model and hardware setup. This initial profiling time is negligible as only a single batch of requests needs to be run on the GPU.
Figure 20 shows the coefficient of determination ($R^{2}$ values) when estimating waiting times for increasing queue sizes across the different models. Overall, we confirm our observation that as the number of request groups in the queue increase the waiting time estimation becomes more accurate. Specifically, we find with four request groups the RWT estimator reaches an accuracy of 0.99.
While the RWT estimator is highly accurate in estimating request waiting time for longer request queues, it is not perfect. When request queues are small, statistical averaging effects of continuous batching do not hold and the waiting time estimation tends calculates more conservative (i.e. higher) estimates leading to lower estimation accuracy.
@@
Impact of Request Group Size. QLM sets the request group size as a multiple ($\delta$) of the average batch size. The exact $\delta$ value depends on the acceptable trade-off between the overhead of running the global scheduler and the granularity of decision-making. As $\delta$ becomes smaller, QLM achieves a finer granularity of decision-making, leading to improved performance. However, the overhead leads to delayed decision-making. Figure 20 demonstrates this tradeoff between performance degradation (caused by changing granularity in decision making) and overhead of the global scheduler when varying $\delta$. At $\delta=16$, the overhead is smallest, but decision-making granularity is coarse, leading to sub-optimal decisions (such as imbalance between virtual queue sizes of LLM serving instances). In contrast, at $\delta=1$, the performance degradation is minimal, but overhead is much higher. We choose $\delta=4$, as it results in nearly zero performance degradation, compared to $\delta=1$, while maintaining a low overhead.
Scalability and Overhead. The global scheduler is invoked only when QLM detects an SLO violation using the RWT estimator. While the global scheduler is performing the virtual queue reordering, QLM continues serving and requests with tight SLOs (low latency requirements) placed at the front of the virtual queues are not interrupted. This ensures the global scheduler is off the critical serving path (i.e. the overheads can be hidden) and ensures SLO compliance. While the global scheduler is the primary overhead, other system overheads include the CPU memory required for the request eviction and model swapping LSOs. Specifically, we require an additional 80 GB CPU memory for Vicuna 13B and Mistral 7B, and 320 GB CPU memory for Llama 70B.
@@
In Figure 20, we show the time required to solve for the linear program in the global scheduler with varying queue sizes in terms of the number of requests. As the basic unit of the solver is a single request group, the model and GPU configurations with a larger request group size would be able to handle a much larger queue size for the same overhead. Consequently, configurations with a large request group size, such as an A100 with a 7B model, can handle a maximum queue size of 400K requests at a 5s overhead per request group (i.e., 5 ms per request, a production requirement).
9. Discussion and Future Work
What are the failure scenarios for the RWT Estimator? The RWT Estimator has the following drawbacks: (a) If the number of request groups in the queue are small, QLM overestimates the request waiting time. Such a conservative estimate is important as it would still lead to SLO preservation by the global scheduler, albeit the system could be underutilized compared to optimal. (b) When the number of output tokens is out-of-distribution (for example, more than 5-10 standard deviations above mean), the estimator would underestimate the request group completion time leading to potential SLO violations. However, in our experience with shareGPT and production traces such scenarios are rare.
What happens if QLM is unable to meet SLOs? QLM’s global scheduler may not be able to find an optimal virtual queue ordering if the request demand is high and the number of LLM serving instances (i.e., underlying compute resources) is insufficient. In such cases, we have three choices: (a) scale up the number of LLM serving instances by adding GPU devices, as we demonstrate in Figure 14 and Figure 11, (b) fall back to a heuristic such as Earliest Deadline First (EDF) and continue serving requests, and (c) performance admission control or rate limiting by dropping incoming requests to limit queue size. Option (a) can only be used when there are available resources, whereas Option (b) and Option (c) would lead to SLO violations.
Can new LSOs be added to QLM? QLM can be extended to support other LSOs that depend on the queue size and request ordering. For example, GPU partitioning techniques (such as NVIDIA MIG (NVIDIA, 2024a)) can be added as an LSO with additional constraints on memory in the linear programming solver described in section 7. We leave the addition of extra LSOs to our future work.
How can QLM handle request priorities? QLM can also be used when strict priorities are assigned to requests instead of SLO values. In the strict priority model, $request_{1}$ would execute before $request_{2}$, if $priority(request_{1})$ is less than $priority(request_{2})$. With strict priority, the relative ordering of requests across priorities is pre-decided, however per-priority assignment still needs to be optimized to maximize SLO satisfaction. Consequently, the concepts of virtual queues, request groups, and the RWT estimator continue to remain useful.
Can SLOs be defined on end-to-end latency? QLM addresses the problem of optimizing time to first token (TTFT) SLO attainment. However, it can also be made to work with the modified objective of end to end request completion time SLOs (i.e. including decoding iterations). In this case, the RWT estimator and global scheduler need to be modified to consider a distribution of number of output tokens per request (as the exact number are unknown aprioiri).
Can SLOs be defined on inter-token latency? QLM primarily addresses the problem of TTFT and does not guarantee the inter-token latency (ITL). However, related work such as Andes (Liu et al., 2024) attempts to guarantee token generation speed (that can be converted into inter-token latency) by modifying the local vLLM scheduler. QLM can work together with such systems to ensure ITL SLOs along with TTFT SLOs.
10. Related Work
  Category Multi Instance SLO-aware Autoregressive Optimizations Hardware Heterogeneity
  LLM Scheduling Optimizations (Yu et al., 2022; Kwon et al., 2023; HuggingFace, 2024)
LLM Serving Backend Optimizations (Liu et al., 2023b; Sheng et al., 2023b; Lienhart, 2024; Zhang et al., 2023a; Liu et al., 2023a; Ge et al., 2024; Miao et al., 2023; Abhyankar et al., 2024; Zhong et al., 2024; Patel et al., 2023; Sheng et al., 2023a; Zhu et al., 2024; Fang et al., 2021; Leviathan et al., 2023)
General ML Model-Serving Systems (Crankshaw et al., 2017; Zhang et al., 2019; Crankshaw et al., 2020; Zhang et al., 2023b; Gujarati et al., 2020)
LLM Orchestration Systems (Sun et al., 2024)
  QLM
  • ✓: Supported, ✗: Not Supported
Table 3. Comparison of QLM with Related Work